OpenStudy (anonymous):

e^2x+3e^x-10=0 Solve the exponential equation. Express the solution as an exact answer in terms of natural logarithms. Then use a calculator to give an approximation of the solution, rounded to two decimal places.

zepdrix (zepdrix):

Well hello there Miss Britt c: \[\Large\rm e^{2x}+3e^x-10=0\]

zepdrix (zepdrix):

Recall one of your exponent rules:\[\Large\rm (a^b)^c=a^{bc}\]Example:\[\Large\rm (x^3)^2=x^{3\cdot2}=x^6\]

zepdrix (zepdrix):

We want to use this in reverse:\[\Large\rm e^{2\cdot x}=(e^x)^2\]

zepdrix (zepdrix):

So we have:\[\Large\rm (e^x)^2+3(e^x)-10=0\]

zepdrix (zepdrix):

What happens if we let \(\Large\rm u=e^x\) or something like that? We get,\[\Large\rm (u)^2+3(u)-10=0\]\[\Large\rm u^2+3u-10=0\]

zepdrix (zepdrix):

Ooo so it's really just a quadratic in disguise! Maybe it factors! Does it? :O hmmm you tell me..

OpenStudy (shadowlegendx):

mhm

OpenStudy (anonymous):

(u+5)(u-2) Sorry I fell asleep last night ha :(

zepdrix (zepdrix):

\[\Large\rm \left(\color{orangered}{u}+5\right)\left(\color{orangered}{u}-2\right)=0\]Which leads to\[\Large\rm u=-5\]\[\Large\rm u=2\]Undoing our substitution,\[\Large\rm e^x=-5\]\[\Large\rm e^x=2\]Take the natural log of each side to solve for x. One of these solutions should end up being extraneous.